# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multi-modal Foundation Models

### v4

@@ 4.2.4 From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables. This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
@@
$p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\alpha}){\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\neq}p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\beta}),\forall{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\in}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G}{\color[rgb]{1,.5,0}{}_{k}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\alpha\neq\beta}.\vspace{-2mm}$
@@
$p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{j}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\alpha}){\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}=}p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{j}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\beta}),\forall{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{j}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\notin}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G}{\color[rgb]{1,.5,0}{}_{k}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\alpha\neq\beta}.\vspace{-2mm}$
@@
Algorithm 1 Generate Prompt from Inductive Bias by LatentExplainer
1:Inductive Bias Formula(s) $\mathcal{F}$, Few-Shot Examples $\mathcal{H}$, pre-trained LLM $\pi_{\theta}$, an optional input of the information about the symbol $\mathcal{I}$.
2:Generated Prompt $\mathcal{P}$
3:function $\mathcal{F}$, $\mathcal{H}$, $\pi_{\theta}$)
@@
7:         semantic $\leftarrow$ $\pi_{\theta}([{\text{symbol},\mathcal{H},\mathcal{I}}])$
@@
10:     $\mathcal{P}=\pi_{\theta}([\mathcal{F},\mathcal{H},\text{semantics}])$
11:     return $\mathcal{P}$
12:end function
@@ 4.3 Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ using in-context learning. The pseudo-code of this section can be found at Algorithm 1 and a more detailed example is at Appendix  C.
Our approach starts by extracting mathematical symbols from the given formula $\mathcal{F}$ using the ExtractSymbols function (line 2). This function traverses the formula to identify and isolate the mathematical symbols.
Next, the algorithm initializes an empty dictionary semantics to store the semantic representations of these symbols (line 3). For each symbol, the pre-trained LLM $\pi_{\theta}$ extracts its semantic meaning based on few-shot examples $\mathcal{H}$ and the optional input symbol information $\mathcal{I}$ (lines 4-6).
Finally, the algorithm generates the prompt $\mathcal{P}$ using the formula $\mathcal{F}$, the few-shot examples $\mathcal{H}$, and the gathered semantics (line 8). This step-by-step reasoning process ensures that the generated prompts are contextually relevant and aligned with the underlying formulas, which could reduce hallucination and enhance model performance.
@@ 4.4 Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses Lin et al. (2023) and image explanations Zhao et al. (2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. The implementation details is in Appendix B.
@@ 5 Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We use GPT-4o as our MLLM and set $temperature=1$ and $top\_p=1.$ The code is available at  https://anonymous.4open.science/r/LatentExplainer-89AF.
@@ 5.1 Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (Karras et al., 2018), AFHQ (Choi et al., 2020), LSUN-Church (Yu et al., 2015) for the unconditional and conditional diffusion models, and CelebA-HQ (Karras et al., 2018), 3DShapes (Burgess and Kim, 2018), and dSprites (Matthey et al., 2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
DDPM GPT-4o Baseline 2.90 13.81 11.84 7.08 87.33 -3.32
+ LatentExplainer w/o IB 3.70 14.22 12.17 7.90 87.30 -3.31
+ LatentExplainer w/o UQ 16.81 31.99 29.04 19.79 89.88 -3.08
+ LatentExplainer 25.11 39.11 36.89 25.97 91.19 -2.98
$\beta$-TCVAE GPT-4o Baseline 5.53 27.70 19.43 10.73 88.97 -3.14
+ LatentExplainer w/o IB 3.85 27.71 20.68 10.55 88.93 -3.15
+ LatentExplainer w/o UQ 18.70 41.13 33.84 22.63 91.03 -2.86
+ LatentExplainer 22.63 47.35 35.63 24.88 91.91 -2.80
Table 2: Quantitative explanation comparison of the disentanglement bias.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
CSVAE GPT-4o Baseline 5.79 28.03 20.24 10.58 89.05 -3.02
+ LatentExplainer w/o IB 8.73 27.32 21.25 12.18 89.16 -3.03
+ LatentExplainer w/o UQ 16.21 35.60 27.58 22.81 90.57 -2.80
+ LatentExplainer 25.85 41.54 35.22 28.51 91.29 -2.74
Table 3: Quantitative explanation comparison of the combination bias.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
Stable Diffusion GPT-4o Baseline 7.07 25.65 21.79 14.72 89.70 -3.27
+ LatentExplainer w/o IB 8.35 26.32 22.94 14.77 89.61 -3.25
+ LatentExplainer w/o UQ 14.85 32.83 27.03 22.02 90.61 -3.12
+ LatentExplainer 16.70 36.32 29.97 23.00 90.97 -3.08
CSVAE GPT-4o Baseline 6.48 22.98 17.03 8.55 88.49 -3.11
+ LatentExplainer w/o IB 9.21 26.68 21.09 11.04 88.98 -3.05
+ LatentExplainer w/o UQ 11.82 29.56 22.07 14.20 88.89 -3.01
+ LatentExplainer 23.70 39.29 32.16 20.33 90.53 -2.89
Table 4: Quantitative explanation comparison of the conditional bias.
@@ 5.2 Generative Models under Inductive Biases
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE (Chen et al., 2018) explicitly penalizes the total correlation of the latent variables to disentangle the latent representations. Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE (Klys et al., 2018) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion (Rombach et al., 2022) is a latent diffusion model to generate images conditioned on prompts.
5.3 Human Annotaions
The ground-truth annotations of explanations are performed by four annotators from the United States and China. All annotators are students that had at least an undergraduate degree. Annotators were presented with the same images and the prompts as MLLM and were asked to annotate the pattern of the images. If there is no clear pattern, just write "No clear explanation". The annotations are then aggregated as references to calculate the automated evaluation metrics.
5.4 Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), SPICE Anderson et al. (2016), BERTScoreZhang et al. (2019), and BARTScore Yuan et al. (2021) as the automated metrics to assess the generated explanations. BLEU, ROUGE-L, METEOR are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive quantitative analysis across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we compare the removal of inductive bias prompts (IB), the removal of uncertainty quantification (UQ), and the full model against the baseline GPT-4o. The results are weighted means across three datasets in Tables 2, Tables 3, and Tables 4. The specific results for each dataset are provided in Table 5 and Table 6 in the Appendix. Removing inductive bias prompts leads to a substantial drop in all generative models. Their consistent results demonstrate that inductive bias is the most important and necessary component when explaining the latent representations of generative models. The removal of uncertainty quantification also results in a slightly decreased performance in all generative models, indicating that uncertainty quantification is also effective, though not as critical as inductive bias prompts. The full model, which incorporates both inductive bias prompts and uncertainty quantification, achieves the highest overall performance, and outperforms all baselines across all models. This confirms the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models. It also shows that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy of latent explanations.

### v5

@@ 4.2.4 From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables Wang et al. (2024). This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
@@
$p({\color[rgb]{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}p_{k}=\alpha}){\color[rgb]{0,1,1}\neq}p({\color[rgb]{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}p_{k}=\beta}),\forall{\color[rgb]{.75,0,.25}z_{i}}{\color[rgb]{0,0,1}\in}{\color[rgb]{1,0,1}G}{\color[rgb]{1,.5,0}{}_{k}},{\color[rgb]{.75,.5,.25}\alpha\neq\beta}.$
@@
Formula:
$p({\color[rgb]{.75,0,.25}z_{j}}\mid{\color[rgb]{.75,.5,.25}p_{k}=\alpha}){\color[rgb]{0,.5,.5}=}p({\color[rgb]{.75,0,.25}z_{j}}\mid{\color[rgb]{.75,.5,.25}p_{k}=\beta}),\forall{\color[rgb]{.75,0,.25}z_{j}}{\color[rgb]{0,0,1}\notin}{\color[rgb]{1,0,1}G}{\color[rgb]{1,.5,0}{}_{k}},{\color[rgb]{.75,.5,.25}\alpha\neq\beta}.$
@@
Algorithm 1 The LatentExplainer Algorithm
1:Input: Inductive Bias Formula(s) $\mathcal{F}$, Optional Information About the Symbol $\mathcal{I}$
2:Require: Few-Shot Examples $\mathcal{H}$, Pre-trained LLM $\pi_{\theta}$, Pre-trained MLLM $\pi_{\theta^{\prime}}$, Generative Model Decoder $\mathcal{G}$
3:Output: Final Explanation $\hat{r}$
@@
7:     semantic $\leftarrow$ $\pi_{\theta}({\text{symbol},\mathcal{H},\mathcal{I}})$
@@
10:$\mathcal{P}=\pi_{\theta}(\mathcal{F},\mathcal{H},\text{semantics})$triangleright Generate inductive bias prompt
11:$\tilde{z}\leftarrow\textsc{ManipulateLatent}(z,\mathcal{F},\mathcal{P})$triangleright Latent variable perturbation
12:$D_{i}\leftarrow\mathcal{G}(\tilde{z})$triangleright Generate image sequence
$\mathcal{R}\leftarrow\emptyset$triangleright Initialize explanation set
14:for $i=1$ to $n$ do
15:     $r_{i}\leftarrow\pi_{\theta^{\prime}}(D_{i},\mathcal{P})$
16:     $\mathcal{R}\leftarrow\mathcal{R}\cup\{r_{i}\}$
17:end for
18:$s_{i}=\frac{1}{n-1}\sum_{i\neq j}\textsc{Sim}(r_{i},r_{j})$
19:$\overline{s}=\frac{1}{n\cdot(n-1)}{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\textsc{Sim}(r_{i},r_{j})$triangleright Compute certainty score
20:$\hat{r}\leftarrow\arg\max_{r_{i}\in\mathcal{R}}s_{i}$
21:if $\overline{s}\geq\epsilon$ then
22:     return $\hat{r}$
23:else
24:     return “No clear explanation”
25:end if
@@ 4.3 Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three bias using in-context learning. A more detailed example is at Appendix D.
As Algorithm 1 shows, our approach starts by extracting mathematical symbols from the given formula $\mathcal{F}$ using the ExtractSymbols function (line 4). This function traverses the formula to identify the mathematical symbols.
Next, the algorithm initializes an empty dictionary semantics to store the semantic representations of these symbols (line 5). For each symbol, the pre-trained LLM $\pi_{\theta}$ extracts its semantic meaning based on few-shot examples $\mathcal{H}$ and the optional input symbol information $\mathcal{I}$ (lines 6-9).
Finally, the algorithm generates the prompt $\mathcal{P}$ using the formula $\mathcal{F}$, the few-shot examples $\mathcal{H}$, and the gathered semantics (line 10). This step-by-step reasoning process ensures that the generated prompts are contextually relevant and aligned with the underlying formulas, which could reduce hallucination and enhance model performance.
4.4 Inductive-bias-guided Data Manipulation
First, identify the relevant formulas within the input inductive bias formulas with regard to a specific latent variable $z_{i}$ to be explained. Then, combine the identified relevant formulas with the inductive bias prompt $\mathcal{P}$ obtained from Section 4.3 and utilize an LLM as a coding agent to generate prompts that specify the modifications needed in the generative model’s decoder code (e.g., adjusting indices through the subscripts of the latent variables or adding properties according to extra information about the symbol $\mathcal{I}$). The coding agent then tests and refines the generated code to effectively perturb $z_{i}$. Through a series of perturbations on $z_{i}$, a sequence of generated images can be obtained, capturing the variations in $z_{i}$ and reflecting its influence on the generated data. Finally, all image sequences generated from the relevant formulas are aggregated to generative explanations about $z_{i}$. The implementation details are in Appendix A.
@@ 4.5 Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses Lin et al. (2023) and image explanations Zhao et al. (2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. Our final explanation $\hat{r}$ is the response that has the highest pairwise cosine similarity with other responses if the latent variable is interpretable. Otherwise, $\hat{r}$ will be “No clear explanation”. The implementation details are in Appendix B.
@@ 5 Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We use GPT-4o as our MLLM backbone set $temperature=1$ and $top\_p=1.$ The code is available at  https://anonymous.4open.science/r/LatentExplainer-89AF.
@@ 5.1 Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ Karras et al. (2018), AFHQ Choi et al. (2020), LSUN-Church Yu et al. (2015) for the unconditional and conditional diffusion models, and CelebA-HQ Karras et al. (2018), 3DShapes Burgess and Kim (2018), and dSprites Matthey et al. (2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
Model Method AFHQ LSUN-Church CelebA-HQ
[…]
DDPM (Disentanglement Bias) Gemini 1.5 Pro 4.28 27.03 11.94 88.68 -3.33 5.38 24.35 8.88 87.45 -3.46 4.94 25.28 8.96 88.82 -3.08
Claude 3.5 Sonnet 12.98 28.7 18.68 89.24 -3.12 12.99 31.71 18.75 89.50 -3.25 13.17 31.62 14.99 89.51 -2.95
GPT-4o 3.08 14.83 8.78 87.55 -3.39 5.52 23.06 11.23 88.47 -3.41 0.11 3.54 1.22 85.97 -3.16
+ LatentExplainer w/o IB 2.14 11.49 7.82 87.17 -3.41 8.78 27.01 14.08 88.85 -3.37 0.19 4.17 1.80 85.88 -3.16
+ LatentExplainer w/o UQ 16.12 31.57 21.70 89.99 -3.12 21.19 37.03 23.65 90.49 -3.19 13.11 27.38 14.03 89.15 -2.92
+ LatentExplainer 25.91 37.84 29.87 91.48 -2.97 30.92 44.21 29.23 91.80 -3.06 18.49 35.27 18.81 90.30 -2.90
Stable Diffusion (Conditional Bias) Gemini 1.5 Pro 0.00 20.52 12.99 88.87 -3.31 2.29 21.53 8.15 88.08 -3.53 0.00 21.46 7.51 88.43 -3.11
Claude 3.5 Sonnet 6.43 29.90 17.54 89.59 -3.24 0.00 28.53 13.8 88.85 -3.47 4.22 25.71 13.63 89.20 -3.10
GPT-4o 7.61 26.32 17.19 90.04 -3.32 7.80 26.73 14.28 89.44 -3.46 5.79 23.89 12.68 89.63 -3.04
+ LatentExplainer w/o IB 9.12 26.59 16.28 89.77 -3.29 10.07 28.31 15.85 89.92 -3.42 5.87 24.06 12.17 89.15 -3.05
+ LatentExplainer w/o UQ 12.36 27.73 20.01 90.03 -3.18 17.19 35.33 21.76 90.73 -3.31 15.01 35.43 23.99 91.07 -2.87
+ LatentExplainer 13.27 29.98 20.08 90.17 -3.19 18.33 38.14 25.01 90.99 -3.23 18.50 40.85 23.90 91.75 -2.82
Table 2: Quantitative results for diffusion models across datasets. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
Model Method 3DShapes CelebA-HQ dSprites
[…]
$\beta$-TCVAE (Disentanglement Bias) Gemini 1.5 Pro 9.31 31.31 8.63 89.50 -3.12 4.56 35.10 12.82 89.87 -3.11 0.98 16.47 7.67 86.39 -3.25
Claude 3.5 Sonnet 10.37 31.39 13.3 89.13 -2.94 11.81 32.89 13.08 89.92 -3.01 5.45 29.72 15.54 89.91 -2.97
GPT-4o 5.51 29.77 7.92 89.18 -3.13 11.09 37.14 17.24 90.55 -3.07 0.00 16.19 7.04 87.17 -3.22
+ LatentExplainer w/o IB 5.41 31.66 10.07 89.44 -3.13 6.14 32.73 14.04 89.92 -3.11 0.00 18.73 7.54 87.42 -3.20
+ LatentExplainer w/o UQ 16.99 37.37 22.89 90.83 -2.80 21.95 48.84 22.60 92.09 -2.88 17.16 37.19 22.40 90.18 -2.89
+ LatentExplainer 25.40 49.06 22.78 91.90 -2.75 29.93 55.68 30.17 93.55 -2.77 12.55 37.30 21.69 90.28 -2.87
CSVAE (Combination Bias) Gemini 1.5 Pro 6.86 25.14 5.88 89.08 -2.98 8.41 24.84 12.49 89.36 -3.09 0.00 16.86 5.6 87.17 -3.21
Claude 3.5 Sonnet 16.11 28.98 21.05 89.05 -2.81 17.98 28.61 20.73 89.50 -2.95 14.28 27.58 18.98 $8$2.94
GPT-4o 6.93 25.07 9.70 89.42 -2.87 10.44 37.42 15.06 90.53 -3.01 0.00 21.61 6.99 87.19 -3.17
+ LatentExplainer w/o IB 14.22 28.57 11.79 90.02 -2.85 11.96 37.02 17.49 90.50 -3.01 0.00 16.36 7.26 86.96 -3.24
+ LatentExplainer w/o UQ 34.55 39.62 30.30 90.83 -2.62 14.08 38.68 22.49 91.11 -2.87 0.00 28.49 15.65 89.78 -2.91
+ LatentExplainer 36.18 43.58 36.75 91.72 -2.52 25.34 45.18 26.96 92.09 -2.81 16.03 35.85 21.82 90.05 -2.90
CSVAE (Conditional Bias) Gemini 1.5 Pro 9.94 28.88 8.06 89.36 -3.10 2.42 17.96 7.02 88.02 -3.10 3.64 25.51 10.89 89.27 -3.10
Claude 3.5 Sonnet 9.00 33.02 15.69 88.48 -2.93 11.75 32.45 19.10 90.68 -2.93 0.00 26.21 16.88 88.62 -3.07
GPT-4o 10.97 30.55 8.83 90.10 -3.02 8.46 28.78 15.78 89.62 -2.98 0.00 9.62 1.03 85.74 -3.34
+ LatentExplainer w/o IB 19.28 39.90 15.74 90.87 -2.88 8.35 28.75 12.11 89.60 -2.97 0.00 11.38 5.26 86.47 -3.29
+ LatentExplainer w/o UQ 16.73 32.28 16.19 89.84 -2.86 13.35 37.20 20.05 89.89 -2.92 5.39 19.20 6.36 86.94 -3.26
+ LatentExplainer 25.71 40.00 20.67 90.88 -2.79 24.21 42.88 19.61 91.21 -2.79 21.17 34.99 20.71 89.49 -3.08
Table 3: Quantitative results for VAE models across datasets. B represents BLEU@4, R represents ROUGE-L, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
5.2 Models and Baselines
Our evaluation benchmarks our proposed LatentExplainer framework against three state-of-the-art multimodal models with strong vision-language reasoning capabilities: GPT-4o OpenAI (2024), Gemini 1.5 Pro Team et al. (2024), and Claude 3.5 Sonnet Anthropic (2024). We employ GPT-4o as a zero-shot baseline, comparing it with the addition of the inductive bias prompt, uncertainty quantification, and the full model with both components included.
@@ 5.3 Generative Models under Inductive Biases
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE Chen et al. (2018) explicitly penalizes the total correlation of the latent variables to disentangle the latent representations. Denoising Diffusion Probabilistic Model (DDPM) Ho et al. (2020) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE Klys et al. (2018) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion Rombach et al. (2022) is a latent diffusion model to generate images conditioned on prompts.
@@
5.4 Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU Papineni et al. (2002), ROUGE-L Lin (2004), SPICE Anderson et al. (2016), BERTScoreZhang et al. (2019), and BARTScore Yuan et al. (2021) as the automated metrics to assess the generated explanations. BLEU, and ROUGE-L are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive quantitative analysis across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we compare the removal of inductive bias prompts (IB), the removal of uncertainty quantification (UQ), and the full model against the baseline GPT-4o. The results for each dataset are provided in Table 2 and Table 3. Removing inductive bias prompts leads to a substantial drop in all generative models. Their consistent results demonstrate that inductive bias is the most important and necessary component when explaining the latent representations of generative models. The removal of uncertainty quantification also results in a slightly decreased performance in all generative models, indicating that uncertainty quantification is also effective, though not as critical as inductive bias prompts. The full model, which incorporates both inductive bias prompts and uncertainty quantification, achieves the highest overall performance, and outperforms all baselines across all models. This confirms the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models. It also shows that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy and reduce hallucination.
In addition to GPT-4o, we evaluate the explanation performance of Gemini 1.5 Pro and Claude 3.5 Sonnet across all generative models and datasets. Both Gemini 1.5 Pro and Claude 3.5 Sonnet underperform compared to our LatentExplainer approach. Our method demonstrates substantial gains in all diffusion and VAE models, particularly for BLEU@4, ROUGE-L, and SPICE, suggesting that explicitly verbalizing inductive biases significantly enhances explanation quality. Furthermore, while Claude 3.5 Sonnet achieves high semantic quality, it does not exhibit the same level of consistency across datasets and models as our proposed method. Overall, while GPT-4o, Claude 3.5 Sonnet and Gemini 1.5 Pro provide competitive baselines, our LatentExplainer consistently improves both lexical and semantic quality, demonstrating its effectiveness in enhancing explainability for latent variables.